Express 2i/1+i in the form a+bi, where



a,b∈R.


1
Expert's answer
2022-06-13T14:44:12-0400

Multiply the numerator and denominator by "\overline{1+i}=1-i":

"\frac{2i}{1+i}=\frac{2i(1-i)}{(1+i)(1-i)}=\frac{2i-2i^2}{1^2-i^2}=\frac{2i-2(-1)}{1-(-1)}=\frac{2i+2}{1+1}=\frac{2+2i}{2}=\frac{2}{2}+\frac{2i}{2}=1+i".